# Relational Representation Distillation

### v3

@@ Abstract
Knowledge distillation (KD) is an effective method for transferring knowledge from a large, well-trained teacher model to a smaller, more efficient student model. Despite its success, one of the main challenges in KD is ensuring the efficient transfer of complex knowledge while maintaining the student’s computational efficiency. Unlike previous works that applied contrastive objectives promoting explicit negative instances with little attention to the relationships between them, we introduce Relational Representation Distillation (RRD). Our approach leverages pairwise similarities to explore and reinforce the relationships between the teacher and student models. Inspired by self-supervised learning principles, it uses a relaxed contrastive loss that focuses on similarity rather than exact replication. This method aligns the output distributions of teacher samples in a large memory buffer, improving the robustness and performance of the student model without the need for strict negative instance differentiation. Our approach demonstrates superior performance on CIFAR-100 and ImageNet ILSVRC-2012, outperforming traditional KD and sometimes even outperforms the teacher network when combined with KD. It also transfers successfully to other datasets like Tiny ImageNet and STL-10. Code is available at https://github.com/giakoumoglou/distillers.
@@ 1 Introduction
    We propose a novel relational consistency framework for KD, called Relational Representation Distillation (RRD), that leverages a large memory buffer and introduces a relational consistency loss. This approach aligns the similarity distributions of teacher and student outputs, maintaining relational structures and improving the robustness and generalization of the student model.
@@
    We validate the effectiveness of RRD hthrough comprehensive testing on standard benchmarks, showcasing considerable gains in both accuracy and robustness. RRD surpasses other methods with a 23.68% relative improvement [^1] Average relative improvement is calculated as: $\frac{1}{N}\sum_{i=1}^{N}\frac{\text{Acc}^{i}{\text{RRD}}-\text{Acc}^{i}{\text{KD}}}{\text{Acc}^{i}{\text{KD}}-\text{Acc}^{i}{\text{van}}}$, where $\text{Acc}^{i}{\text{RRD}}$, $\text{Acc}^{i}{\text{KD}}$, and $\text{Acc}^{i}_{\text{van}}$ represent the accuracies of RRD, KD, and vanilla training of the $i$-th student model, respectively [51]. over conventional KD. When integrated with KD, it demonstrates a 68.66% relative improvement over standard KD, underscoring its robust potential in enhancing model performance significantly.
@@
    We provide qualitative evidence through t-SNE visualizations, demonstrating that RRD effectively preserves the spatial relationships in the embedding spaces of both student and teacher models, leading to improved consistency in feature representations and enhanced student performance.
@@ 2 Related Work
Self-supervised learning. Self-supervised learning has significantly impacted representation learning by leveraging unlabeled data. Various pretext tasks have been proposed to learn these representations such as colorizing an image [61], predicting a missing patch [43], estimating the rotation angle [16], jigsaw puzzle [39, 40]. Contrastive methods further extend this concept by pulling together an anchor and a "positive" sample in the embedding space while pushing apart the anchor from many "negative" samples [26], represented by frameworks like SimCLR [8], MoCo [18, 9, 10] and other variants [56, 25]. Different from these, ReSSL [63] introduces relational self-supervised learning, which explores the relationships between data points. These approaches have inspired various KD methods, including our proposed method, which adapts relational consistency from self-supervised learning to the KD framework.
Knowledge distillation. The study of [22] on KD introduced the idea of transferring knowledge from large, complex models to smaller, more efficient models without losing their generalization capabilities. This technique involves temperature scaling in the softmax outputs to effectively capture and convey the knowledge from the teacher model. Subsequent methods have largely focused on enhancing the KD objective by incorporating additional loss functions, often complementing the standard cross-entropy and KL-divergence losses. For example, using intermediate representations or "hints" to guide the learning process [46], and aligning the attention maps of teacher and student models [59]. Other methods preserve relational knowledge between samples [52] or align the correlation structures between teacher and student models [44]. Additionally, some techniques use variational inference to improve the knowledge transfer process [1]. Further developments include focusing on structural relationships between data points to ensure the student model learns relational information [41], and maintaining the internal dynamics of neural networks during distillation [42, 21]. Other notable techniques include network compression through factor transfer [27], optimizing networks for better transfer learning [58], promoting selectivity in the distillation process [24]. Building upon the influence of self-supervised learning, CRD [51] is the first study that combines contrastive learning with KD, which aims to maximize mutual information between the teacher and student representations. CRD utilizes a memory buffer [55], which allows it to maintain a large set of negative samples. This approach enables CRD to perform contrastive learning effectively in the KD setting, enhancing the quality of transferred knowledge. Other approaches use cross-entropy with label smoothing [64], estimate propensity scores to compensate for under-represented samples [38], or apply $\ell_{2}$ loss to align teacher and student features [57]. Some techniques preserve relations between predictions [23], employ orthogonal matrices for intra-batch feature similarity [34], or automatically assign proper target layers [4]. Researchers have introduced cross-stage connection paths [7] and direct reuse of the teacher’s classifier [3]. They have also explored differentiable meta-learning [14] and many-to-one representation matching mechanisms [30]. Studies have leveraged complementary losses to maximize correlation and mutual information [36], pairwise feature kernel comparisons [17], and second-order gradient information [65]. The field has incorporated graph-based methods [33] and contrastive intra-batch losses [6]. Some approaches derive supervision from the student’s intermediate features without additional teacher models [2], while others propose training-free frameworks for architecture search [15]. Additional work focuses on normalizing logits before applying softmax and KL divergence [50].
@@ 3 Methodology
This section presents our methodology to improve the efficiency and accuracy of KD. Our method, Relational Representation Distillation (RRD), focuses on maintaining relational consistency between the teacher and student models by leveraging a large memory buffer of teacher samples to align their output distributions. By ensuring consistent relational structures, RRD enhances the robustness and performance of the student model. Figure 1 shows an overview of the proposed RRD method.
Figure 1: Overview of the RRD method. The student network processes input $x_{i}$ to produce embeddings $z_{i}^{S}$, while the teacher network, indicated as frozen by a snowflake, processes the same input to generate embeddings $z_{i}^{T}$. The embeddings are stored in a memory buffer to align their output distributions via softmax layers. The relational consistency between the teacher and student models is enforced by leveraging the stored embeddings to train the student network.
@@ 3.1 Preliminaries
KD involves transferring knowledge from a high-capacity teacher neural network, denoted as $f^{T}$, to a more compact student neural network, $f^{S}$. Consider $x_{i}$, with $i=1,2,\ldots,N$, as the input to these networks, typically an image. We represent the outputs at the penultimate layer (just before the final classification layer, or logits) as $z_{i}^{T}=f^{T}(x_{i})$ and $z_{i}^{S}=f^{S}(x_{i})$ for the teacher and student models, respectively. The primary objective of KD is to enable the student model to approximate the performance of the teacher model while leveraging the student’s computational efficiency. The overall distillation process can be mathematically expressed as:
$\mathcal{L}=\mathcal{L}_{\text{sup}}(y_{i},z_{i}^{S})+\lambda\cdot\mathcal{L}_{\text{distill}}(z_{i}^{T},z_{i}^{S})$ (1)
@@ 3.2 Relational Consistency
Our method is inspired by self-supervised learning. Traditional contrastive self-supervised learning uses instance discrimination as a pretext task and relies on ($K+1$)-softmax classification, where different instances are pushed apart, and augmented views of the same instance are expected to have identical features. Applying such properties to KD imposes overly strict constraints, where a contrastive loss encourages the representations from the teacher and student models for the same input data to be similar, while simultaneously pushing apart representations from different data inputs:
$\mathcal{L}_{\text{contrast}}(z_{i}^{T},z_{i}^{S})=-\log\frac{\exp(\phi(z_{i}^{T},z_{i}^{S})/\tau)}{\sum_{j=1}^{M}\exp(\phi(z_{i}^{S},z_{j}^{T})/\tau)}$ (2)
where $\mathcal{L}_{\text{contrast}}$ is the InfoNCE [53] loss, typically employed in self-supervised methods, $\phi$ is a similarity function, $\tau$ is a temperature parameter, and $M$ is the number of negative samples.
In this way, we do not encourage explicit negative instances (those to be pushed away) for each instance; instead, we leverage the pairwise similarities to explore their relationships. We pull the features of the student $f^{S}$ and teacher $f^{T}$. As a result, our method relaxes Equation 2, where different teacher and student outputs do not always need to be pushed away from each other; and teacher and student need to share similar but not exactly the same features.
@@
Concretely, given an input image $x_{i}$, and the outputs $z_{i}^{T}=f^{T}(x_{i})$ and $z_{i}^{S}=f^{S}(x_{i})$ for the teacher and student models, respectively, we calculate their similarity measured by $\phi(z_{i}^{T},z_{i}^{S})$. A softmax layer can be adopted to process the calculated similarities, which then produces a relationship distribution:
$p_{i}^{T}=\frac{\exp(\phi(z_{i}^{T},z_{j}^{T})/\tau_{t})}{\sum_{k=1}^{M}\exp(\phi(z_{i}^{T},z_{k}^{T})/\tau_{t})}$ (3)
@@
where $\tau_{t}$ is the temperature parameter for the teacher network. At the same time, we can calculate the relationship distribution for the student model as:
$p_{i}^{S}=\frac{\exp(\phi(z_{i}^{S},z_{j}^{S})/\tau_{s})}{\sum_{k=1}^{M}\exp(\phi(z_{i}^{S},z_{k}^{S})/\tau_{s})}$ (4)
where $\tau_{s}$ is a different temperature parameter for the student network. We propose to push the relational consistency between $p_{i}^{T}$ and $p_{i}^{S}$, similar to [63], by minimizing the KL divergence, which can be formulated as:
$\mathcal{L}_{\text{relational}}(z_{i}^{T},z_{i}^{S})=\textit{KL}(p_{i}^{T}\parallel p_{i}^{S})=H(p_{i}^{T},p_{i}^{S})-H(p_{i}^{T})$ (5)
@@
where KL denotes the KL divergence between $p_{i}^{T}$ and $p_{i}^{S}$. Since $p_{i}^{T}$ will be used as a target, the gradient will be clipped here to avoid the model collapsing, thus we only minimize the cross-entropy term $H(p_{i}^{T},p_{i}^{S})$ in our implementation.
However, the quality of the target similarity distribution $p_{i}^{T}$ is crucial for reliable and stable training. To ensure this, we maintain a large memory buffer $Q$, storing the feature embeddings from teacher batches. The relational consistency between the teacher and student models is enforced by aligning the similarity distributions of their outputs using the KL divergence.
We normalize the outputs $z_{i}^{T}$ and $z_{i}^{S}$ before computing the loss, ensuring that the representations lie on a unit hypersphere. To compute $\mathcal{L}_{\text{relational}}$, we further encode $z_{i}^{T}$ and $z_{i}^{S}$ using a projection head to match the dimensions. This ensures that the representations from both models are compatible for comparison and alignment.
3.3 Final Objective
@@
$\mathcal{L}=\mathcal{L}_{\text{sup}}(y_{i},z_{i}^{S})+\lambda\cdot\mathcal{L}_{\text{distill}}(z_{i}^{T},z_{i}^{S})+\beta\cdot\mathcal{L}_{\text{relational}}(z_{i}^{T},z_{i}^{S})$ (6)
where $\beta$ is a hyperparameter that balances the KD loss $\mathcal{L}_{\text{relational}}$. We experiment with $\beta$ in ablation studies to understand its impact on the final performance.
@@ 4 Experiments
We evaluate our RRD framework in the KD task of model compression of a large network to a smaller one, following the same protocol as [22, 46, 59, 52, 44, 1, 41, 42, 21, 27, 58, 24, 51]. This approach is consistent with the standard practice in the field, allowing for fair comparison across methods.
Datasets. (1) CIFAR-100 [28] contains 50,000 training images with 500 images per class and 10,000 test images. (2) ImageNet ILSVRC-2012 [13] includes 1.2 million images from 1,000 classes for training and 50,000 for validation. (3) STL-10 [12] consists of a training set of 5,000 labeled images from 10 classes and 100,000 unlabeled images, and a test set of 8,000 images. (4) Tiny ImageNet (TIN-200) [13] has 200 classes, each with 500 training images and 50 validaton images.
Setup. We experiment on CIFAR-100 and ImageNet with student-teacher combinations of various capacity, such as ResNet [19] or Wide ResNet (WRN) [60], VGG [49], MobileNet [47], and ShuffleNet [62, 32] (more details about the network architectures are described in the supplementary material). We use $M=16384$ negative samples as in [51] (we ablate $M$ in Section 4.6) and set the temperature parameter of the student to $\tau_{s}=0.04$ and of the teacher to $\tau_{t}=0.07$, also used in [63] that employs dual temperature (we also ablate both temperatures in Section 4.6). Following [51], both the student and teacher outputs are projected to a 128-dimensional space. We use a projection head of a single linear layer, followed by $\ell_{2}$ normalization. We train for a total of 240 epochs. The hyperparameter $\lambda$ is set to $1.0$ for the KL divergence loss to maintain consistency with the baseline CRD method [51], facilitating fair comparison. This value of $\lambda$ is also fixed to $1.0$ in other methods such as [3, 4, 51]. The loss coefficient $\beta$ is set to $1.0$ (we study the impact of $\beta$ in Section 4.6). More details on the training details are described in the supplementary material.
Comparison. We compare our approach to the following state-of-the-art methods: (1) KD [22]; (2) FitNets [46]; (3) AT [59]; (4) SP [52]; (5) CC [44]; (6) VID [1]; (7) RKD [41]; (8) PKT [42]; (9) AB [21]; (10) FT [27]; (11) FSP [58]; (12) NST [24]; (13) CRD [51]; (14) OFD [20]; (15) WSLD [64]; (16) IPWD [38]. In the supplementary material, we include additional methods: (17) SRRL [57]; (18) SemCKD [4]; (19) ReviewKD [7]; (20) SimKD [3]; (21) DistPro [14]; (22) NORM [30]; (23) ITRD [36]; (24) FKD [17]; (25) CRCD [65]; (26) EGA [33]; (27) WCoRD [6].
Role of the projection layer. Following [35], we use a projection head consisting of a single linear layer, followed by $\ell_{2}$ normalization. The projection layer is not merely a dimension-matching tool but plays a crucial role in KD. It effectively transfers knowledge by implicitly encoding relational information from previous samples [35]. The weights within this projection layer capture crucial relational data, which is fundamental for conveying information about the correlation between features, thus significantly influencing the training dynamics.

### v4

@@ Abstract
Knowledge Distillation (KD) is an effective method for transferring knowledge from a large, well-trained teacher model to a smaller, more efficient student model. Despite its success, one of the main challenges in KD is ensuring the efficient transfer of complex knowledge while maintaining the student’s computational efficiency. While contrastive learning methods typically push different instances apart and pull similar ones together, applying such constraints to KD can be too restrictive. Contrastive methods focus on instance-level information, but lack attention to relationships between different instances. We propose Relational Representation Distillation (RRD), which improves knowledge transfer by maintaining structural relationships between feature representations rather than enforcing strict instance-level matching. Specifically, our method employs sharpened distributions of pairwise similarities among different instances as a relation metric, which is utilized to match the feature embeddings of student and teacher models. Our approach demonstrates superior performance on CIFAR-100 and ImageNet ILSVRC-2012, outperforming traditional KD and sometimes even outperforms the teacher network when combined with KD. It also transfers successfully to other datasets like Tiny ImageNet and STL-10. Code is available at https://github.com/giakoumoglou/distillers.
@@ 1 Introduction
    We introduce Relational Representation Distillation (RRD), a KD method that leverages a memory buffer to align the similarity distributions between teacher and student outputs while maintaining structural relationships.
@@
    We validate the effectiveness of RRD through comprehensive testing on standard benchmarks, showcasing considerable gains in both accuracy and robustness. RRD surpasses other methods with a 23.68% relative improvement [^1] Average relative improvement is calculated as: $\frac{1}{N}\sum_{i=1}^{N}\frac{\text{Acc}^{i}_{\text{RRD}}-\text{Acc}^{i}_{\text{KD}}}{\text{Acc}^{i}_{\text{KD}}-\text{Acc}^{i}_{\text{van}}}$, where $\text{Acc}^{i}_{\text{RRD}}$, $\text{Acc}^{i}_{\text{KD}}$, and $\text{Acc}^{i}_{\text{van}}$ represent the accuracies of RRD, KD, and vanilla training of the $i$-th student model, respectively [46]. over conventional KD. When integrated with KD, it demonstrates a 68.66% relative improvement over standard KD.
@@
    We provide qualitative evidence through t-SNE visualizations, demonstrating that RRD effectively preserves the spatial relationships in the embedding spaces of both student and teacher models,
@@ 2 Related Work
Self-supervised learning has significantly impacted representation learning by leveraging unlabeled data. Various pretext tasks have been proposed to learn these representations, such as colorizing an image [56], predicting a missing patch [39], estimating the rotation angle [14], and solving jigsaw puzzles [35, 36]. Contrastive methods further extend this concept by pulling together an anchor and a ”positive” sample in the embedding space while pushing apart the anchor from many ”negative” samples [24], represented by frameworks such as [7, 17, 8, 9] and other variants [51, 23]. Differing from these, [58] introduces relational self-supervised learning, which explores the relationships between data points. These approaches have inspired various KD methods, including our proposed method, which adapts relational consistency from self-supervised learning to the KD framework.
Knowledge distillation.
The work by [20] laid the foundation for knowledge distillation (KD) by introducing a method for transferring knowledge from larger teacher models to smaller student models, ensuring that the student retains strong generalization capabilities. This process uses temperature scaling in the teacher’s softmax output to produce ”soft targets” that are easier for the student to learn from. Many methods extend this approach by adding loss functions that complement standard cross-entropy and KL-divergence losses. Examples include utilizing intermediate representations or ”hints” to guide student learning [42], aligning attention maps between teacher and student [54], and preserving relational information between samples [47].
Feature-based knowledge distillation.
Feature-based KD methods focus on transferring knowledge through intermediate representations or feature alignments. These approaches often involve aligning the spatial and channel-wise attention maps of teacher and student models, enabling the student to better mimic the teacher’s feature extraction [54]. Other techniques directly align the teacher and student’s correlation structures [40], or employ variational inference to optimize knowledge transfer [1]. Additionally, relational approaches maintain structural relationships in the feature space, ensuring the student model captures essential relational knowledge [37]. Methods preserving internal dynamics across layers have also shown to be effective in sustaining feature-based information during KD [38, 19].
Contrastive knowledge distillation.
Contrastive KD methods combine self-supervised contrastive learning with KD principles to improve knowledge transfer. A notable approach is CRD [46], which applies contrastive loss to maximize mutual information between teacher and student representations. CRD uses a memory buffer [50] to maintain a large set of negative samples, which enables effective contrastive learning and enhances the quality of knowledge transfer. These contrastive approaches aim to preserve the teacher’s feature space, resulting in better student model performance in KD tasks.
Relational and structural knowledge distillation.
Several KD techniques focus on preserving structural and relational information in the student model. Methods like relational KD capture relationships between data points, enabling the student model to learn the structural organization of the teacher’s feature space [37]. Techniques like using orthogonal matrices for intra-batch feature similarity [32] and aligning pairwise feature kernels [15] further enhance relational knowledge transfer. This category also includes approaches that optimize relational information among outputs, such as preserving relations between predictions [21], aligning teacher and student features using $\ell_{2}$ loss [52], and maintaining alignment in the student’s prediction space. Cross-stage connections [6] and direct reuse of the teacher’s classifier [3] contribute to better alignment in learned representations, while differentiable meta-learning optimizes layer-wise transfer [13].
@@ 3 Methodology
This section presents our methodology to improve the efficiency and accuracy of KD. Our method, Relational Representation Distillation (RRD), improves KD efficiency by maintaining structural relationships between feature representations using a memory buffer of teacher samples to align their output distributions. Figure 1 shows an overview of the proposed RRD method.
Figure 1: Overview of the RRD method. The student network processes input $x_{i}$ to produce embeddings $\mathbf{z_{i}^{S}}$, while the teacher network, represented as frozen by a snowflake, processes the same input to generate embeddings $\mathbf{z_{i}^{T}}$. The embeddings are stored in a memory buffer to align their output distributions via softmax layers. The relationship between the teacher and student models is reinforced by leveraging the stored embeddings to guide the training of the student network.
@@ 3.1 Preliminaries
KD involves transferring knowledge from a high-capacity teacher neural network, denoted as $f^{T}$, to a more compact student neural network, $f^{S}$. Consider $x_{i}$ as the input to these networks, typically an image. We represent the outputs at the penultimate layer (just before the final classification layer, or logits) as $\mathbf{z_{i}^{T}}=f^{T}(x_{i})$ and $\mathbf{z_{i}^{S}}=f^{S}(x_{i})$ for the teacher and student models, respectively. The primary objective of KD is to enable the student model to approximate the performance of the teacher model while leveraging the student’s computational efficiency. The overall distillation process can be mathematically expressed as:
$\mathcal{L}=\mathcal{L}_{\text{sup}}(y_{i},\mathbf{z_{i}^{S}})+\lambda\cdot\mathcal{L}_{\text{distill}}(\mathbf{z_{i}^{T}},\mathbf{z_{i}^{S}})$ (1)
@@ 3.2 Relational Representation Distillation
Traditional contrastive learning uses instance discrimination and relies on ($K+1$)-softmax classification, where different instances are pushed apart, and matching instances are expected to have identical features. This can lead to the class collision problem where semantically similar instances are forced apart even when they should maintain some similarity. Applying such properties to KD imposes overly strict constraints, where a contrastive loss encourages the representations from the teacher and student models for the same input data to be similar, while simultaneously pushing apart representations from different data inputs:
$\mathcal{L}_{\text{contrastive}}(\mathbf{z_{i}^{T}},\mathbf{z_{i}^{S}})=-\log\frac{\exp(\phi(\mathbf{z_{i}^{T}},\mathbf{z_{i}^{S}})/\tau)}{\sum_{k=1}^{M}\exp(\phi(\mathbf{z_{i}^{S}},\mathbf{z_{k}})/\tau)}$ (2)
where $\phi$ is a similarity function, typically the dot product between $\ell_{2}$ normalized embeddings $\phi(\mathbf{u},\mathbf{v})=\mathbf{u}^{T}\cdot\mathbf{v}/||\mathbf{u}||\cdot||\mathbf{v}||$, $\tau$ is a temperature parameter, and $M$ is the number of negative samples typically sampled from a memory buffer $Q=\{\mathbf{z_{k}}\}_{k=1}^{M}$.
Our method relaxes this contrastive objective by not enforcing strict instance-level matching. Instead of explicitly pushing different instances apart, we focus on preserving pairwise similarity relationships between instances. This relaxation allows the student more flexibility in learning from the teacher where different instances can maintain varying degrees of similarity rather than being strictly separated, and teacher-student pairs do not need to have identical features.
@@
Specifically, to maintain structural relationships between instances, given an input image $x_{i}$, and the outputs $\mathbf{z_{i}^{T}}=f^{T}(x_{i})$ and $\mathbf{z_{i}^{S}}=f^{S}(x_{i})$ for the teacher and student models, respectively, we calculate the similarities between $\mathbf{z_{i}^{T}}$ and the $j$-th instance of the memory bank as $\phi(\mathbf{z_{i}^{T}},\mathbf{z_{j}})$. A softmax layer can be adopted to process the calculated similarities, which then produces a similarity distribution:
$\mathbf{p_{i}^{T}}(j)=\frac{\exp(\phi(\mathbf{z_{i}^{T}},\mathbf{z_{j}})/\tau_{t})}{\sum_{k=1}^{M}\exp(\phi(\mathbf{z_{i}^{T}},\mathbf{z_{k}})/\tau_{t})}$ (3)
@@
where $\tau_{t}$ is the temperature parameter for the teacher network. At the same time, we can calculate the relationship between $\mathbf{z_{i}^{S}}$ and the $j$-th instance of the memory bank as $\phi(\mathbf{z_{i}^{S}},\mathbf{z_{j}})$. The resulting similarity distribution is:
$\mathbf{p_{i}^{S}}(j)=\frac{\exp(\phi(\mathbf{z_{i}^{S}},\mathbf{z_{j}})/\tau_{s})}{\sum_{k=1}^{M}\exp(\phi(\mathbf{z_{i}^{S}},\mathbf{z_{k}})/\tau_{s})}$ (4)
where $\tau_{s}$ is a different temperature parameter for the student network. We set $\tau_{s}<\tau_{t}$ to create sharpened distributions for the student model. That way the student’s similarity distribution $\mathbf{p_{i}^{S}}$ becomes more peaked around the highest similarity values compared to the teacher’s distribution $\mathbf{p_{i}^{T}}$. This sharpening helps the student focus on learning the most significant relationships while maintaining flexibility in capturing secondary similarities, thus avoiding the class collision problem that occurs with strict instance discrimination.
We align these distributions by minimizing their KL divergence, a relation metric to quantify the similarity between the teacher and student distributions:
$\mathcal{L}_{\text{relational}}(\mathbf{z_{i}^{T}},\mathbf{z_{i}^{S}})=D_{\text{KL}}(\mathbf{p_{i}^{T}}\parallel\mathbf{p_{i}^{S}})=H(\mathbf{p_{i}^{T}},\mathbf{p_{i}^{S}})-H(\mathbf{p_{i}^{T}})$ (5)
@@
where $D_{\text{KL}}$ denotes the KL divergence between $\mathbf{p_{i}^{T}}$ and $\mathbf{p_{i}^{S}}$. Since $\mathbf{p_{i}^{T}}$ will be used as a target, the gradient will be clipped here to avoid model collapse, thus we only minimize the cross-entropy term $H(\mathbf{p_{i}^{T}},\mathbf{p_{i}^{S}})$ in our implementation.
The quality of the target similarity distribution $\mathbf{p_{i}^{T}}$ is crucial for reliable and stable training, which we achieve by maintaining a large memory buffer to store feature embeddings from teacher batches. The structural relationships between the teacher and student models are preserved by aligning the similarity distributions of their outputs using the KL divergence. To ensure the representations lie on a unit hypersphere, we normalize the outputs $\mathbf{z_{i}^{T}}$ and $\mathbf{z_{i}^{S}}$ before computing the loss. Furthermore, $\mathcal{L}_{\text{relational}}$ is computed by encoding $\mathbf{z_{i}^{T}}$ and $\mathbf{z_{i}^{S}}$ through a projection head that matches their dimensions, ensuring compatibility for comparison and alignment. This projection head also facilitates knowledge transfer by implicitly encoding relational information from previous samples [30].
@@
$\mathcal{L}=\mathcal{L}_{\text{sup}}(y_{i},\mathbf{z_{i}^{S}})+\lambda\cdot\mathcal{L}_{\text{distill}}(\mathbf{z_{i}^{T}},\mathbf{z_{i}^{S}})+\beta\cdot\mathcal{L}_{\text{relational}}(\mathbf{z_{i}^{T}},\mathbf{z_{i}^{S}})$ (6)
where $\beta$ is a hyperparameter that balances the proposed loss $\mathcal{L}_{\text{relational}}$.
@@ 4 Experiments
We evaluate our RRD framework in the KD task of model compression of a large network to a smaller one [46]. This approach is consistent with the standard practice in the field, allowing for fair comparison across methods.
4.1 Experimental Setup
We implement DCD in PyTorch following the implementation of CRD [46].
Datasets.
(1) CIFAR-100 [26] contains 50,000 training images with 500 images per class and 10,000 test images. (2) ImageNet ILSVRC-2012 [12] includes 1.2 million images from 1,000 classes for training and 50,000 for validation. (3) STL-10 [11] consists of a training set of 5,000 labeled images from 10 classes and 100,000 unlabeled images, and a test set of 8,000 images. (4) Tiny ImageNet [12] has 200 classes, each with 500 training images and 50 validation images.
Setup.
We experiment on CIFAR-100 and ImageNet with student-teacher combinations of various capacity, such as ResNet [16] or Wide ResNet (WRN) [55], VGG [45], MobileNet [43], and ShuffleNet [57, 29] (more details about the network architectures are described in the supplementary material). We use $M=16384$ samples (we ablate $M$ in Section 4.7) and set the temperature parameter of the student to $\tau_{s}=0.04$ and of the teacher to $\tau_{t}=0.07$ (we also ablate both temperatures in Section 4.7). We set $\lambda=1$ and $\beta=1$ (we study the impact of $\beta$ in Section 4.7). Both the student and teacher outputs are projected to a 128-dimensional space. We use a projection head of a single linear layer, followed by $\ell_{2}$ normalization. We train for 240 epochs for CIFAR-100 and 120 for ImageNet. More details on the training details are described in the supplementary material.
Comparison.
We compare our approach to the following state-of-the-art methods: (1) KD [20]; (2) FitNets [42]; (3) AT [54]; (4) SP [47]; (5) CC [40]; (6) VID [1]; (7) RKD [37]; (8) PKT [38]; (9) AB [19]; (10) FT [25]; (11) FSP [53]; (12) NST [22]; (13) CRD [46]; (14) OFD [18]; (15) WSLD [59]; (16) IPWD [34]. In the supplementary material, we include additional methods.
Role of the projection layer.
Following [30], we use a projection head consisting of a single linear layer, followed by $\ell_{2}$ normalization. The projection layer is not merely a dimension-matching tool but plays a crucial role in KD. It effectively transfers knowledge by implicitly encoding relational information from previous samples [30].